from 5 statisticians and 6 mathematicians a committee consists of 3 statistician and 2 mathematicians is to be formed . how many different committees can be formed if:
a. there is no restriction
b. 2 particular statisticians must be in the committee
c. one particular mathematician can not be in the committee

Answer by Boreal(15235)

a. 11C5=462
b. 3C2*6C1=18 ways
c. If one particular mathematician can't be on the committee, the choice is now from 10 people, 5 from each discipline.
5C3*5C2=10*10=100 ways